# Continuous Diffusion for Mixed-Type Tabular Data

### v2

@@ Appendix I Machine Learning Efficiency Models
CoDi is prohibitively expensive to train on lending and diabetes and TabDDPM produces NaNs for acsincome and diabetes. For both models, the performance metrics on these datasets are therefore not reported.
@@
Table 6: Model evaluation results averaged over 11 datasets (skipping a dataset if the model was not trainable on it) for five benchmark models and for CDTD with three different noise schedules. Per performance metric, bold indicates the best, underline the second best result.
@@
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
RMSE (abs. diff.; $\downarrow$) 0.094 0.674 0.947 0.486 0.173 0.084 0.101 0.110
F1 (abs. diff.; $\downarrow$) 0.053 0.130 0.074 0.015 0.044 0.025 0.020 0.025
AUC (abs. diff.; $\downarrow$) 0.020 0.080 0.065 0.009 0.027 0.018 0.016 0.022
L2 distance of corr.  ($\downarrow$) 1.321 2.187 2.745 3.786 1.200 0.782 0.756 0.990
Detection score ($\downarrow$) 0.934 0.986 0.976 0.769 0.936 0.796 0.768 0.783
JSD ($\downarrow$) 0.011 0.114 0.152 0.051 0.038 0.015 0.016 0.018
WD ($\downarrow$) 0.011 0.023 0.025 0.061 0.022 0.010 0.007 0.009
DCR (abs. diff. to test; $\downarrow$) 1.588 3.336 1.621 0.568 1.000 0.796 0.806 0.758
@@
Table 7: L2 norm (incl. standard errors in subscripts) of the correlation matrix differences of real and synthetic train sets for five benchmark models and for CDTD with three different noise schedules.
@@
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.141_{\pm 0.003}$ $0.129_{\pm 0.003}$ $0.119_{\pm 0.002}$
adult $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.156_{\pm 0.006}$ $0.493_{\pm 0.009}$ $0.170_{\pm 0.007}$ $0.125_{\pm 0.009}$ $0.128_{\pm 0.010}$
bank $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.898_{\pm 0.025}$ $0.499_{\pm 0.021}$ $0.323_{\pm 0.008}$ $0.266_{\pm 0.011}$ $0.256_{\pm 0.015}$
beijing $0.133_{\pm 0.006}$ $1.445_{\pm 0.009}$ $1.642_{\pm 0.015}$ $1.133_{\pm 0.035}$ $0.363_{\pm 0.015}$ $0.075_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.071_{\pm 0.004}$
churn $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.327_{\pm 0.044}$ $0.746_{\pm 0.062}$ $0.302_{\pm 0.041}$ $0.289_{\pm 0.043}$ $0.282_{\pm 0.044}$
covertype $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.044_{\pm 0.001}$ $1.029_{\pm 0.032}$ $2.359_{\pm 0.011}$ $2.275_{\pm 0.009}$ $2.710_{\pm 0.009}$
default $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $3.408_{\pm 0.105}$ $1.672_{\pm 0.061}$ $0.627_{\pm 0.068}$ $0.652_{\pm 0.102}$ $0.737_{\pm 0.033}$
diabetes $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $1.201_{\pm 0.020}$ $0.803_{\pm 0.032}$ $1.345_{\pm 0.016}$
lending $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.675_{\pm 0.015}$ $1.042_{\pm 0.075}$ $1.189_{\pm 0.040}$ $1.363_{\pm 0.097}$
news $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $15.985_{\pm 0.081}$ $4.874_{\pm 0.148}$ $1.925_{\pm 0.527}$ $2.035_{\pm 0.475}$ $3.395_{\pm 0.950}$
nmes $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.447_{\pm 0.031}$ $0.609_{\pm 0.032}$ $0.433_{\pm 0.025}$ $0.478_{\pm 0.083}$ $0.481_{\pm 0.058}$
@@
Table 8: Jensen-Shannon divergence (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
@@
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.000}$ $0.022_{\pm 0.001}$
adult $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.034_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.016_{\pm 0.000}$
bank $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.029_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.010_{\pm 0.000}$ $0.009_{\pm 0.001}$ $0.012_{\pm 0.001}$
beijing $0.005_{\pm 0.002}$ $0.147_{\pm 0.003}$ $0.257_{\pm 0.001}$ $0.035_{\pm 0.003}$ $0.018_{\pm 0.004}$ $0.003_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.005_{\pm 0.001}$
churn $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.014_{\pm 0.004}$ $0.043_{\pm 0.001}$ $0.012_{\pm 0.003}$ $0.012_{\pm 0.002}$ $0.011_{\pm 0.002}$
covertype $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.011_{\pm 0.000}$
default $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.027_{\pm 0.002}$ $0.073_{\pm 0.002}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
diabetes $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.023_{\pm 0.000}$ $0.026_{\pm 0.000}$
lending $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.243_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.064_{\pm 0.002}$
news $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.046_{\pm 0.000}$ $0.012_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
nmes $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.028_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.008_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.013_{\pm 0.003}$
@@
Table 9: Wasserstein distance (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
@@
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.007_{\pm 0.000}$ $0.037_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
adult $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
bank $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$
beijing $0.009_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.058_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
churn $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.006_{\pm 0.001}$
covertype $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.015_{\pm 0.000}$ $0.012_{\pm 0.000}$
default $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
diabetes $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.042_{\pm 0.000}$
lending $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.425_{\pm 0.001}$ $0.009_{\pm 0.000}$ $0.010_{\pm 0.000}$ $0.011_{\pm 0.000}$
news $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.078_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
nmes $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$
@@
Table 10: Detection score (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.

### v3

@@ Appendix I Machine Learning Efficiency Models
CoDi is prohibitively expensive to train on lending and diabetes and TabDDPM produces NaNs for acsincome and diabetes. SMOTE takes too long to sample datasets of a sufficient size for acsincome and covertype (see Table 29). For those models, the performance metrics on these datasets are therefore not reported. They are assigned a rank of 10 in Table 1 and are not taken into account when forming the average metrics reported in Table 11.
@@
Table 11: Model evaluation results averaged over 11 datasets (skipping a dataset if the model was not trainable on it, which due to extensive sampling times for SMOTE includes two of the most complex datasets, acsincome and covertype) for seven benchmark models and for CDTD with three different noise schedules. Per performance metric, bold indicates the best, underline the second best result.
@@
SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
RMSE (abs. diff.; $\downarrow$) 0.083 0.094 0.674 0.947 0.486 0.173 0.313 0.084 0.101 0.110
F1 (abs. diff.; $\downarrow$) 0.007 0.053 0.130 0.074 0.015 0.044 0.099 0.025 0.020 0.025
AUC (abs. diff.; $\downarrow$) 0.008 0.020 0.080 0.065 0.009 0.027 0.059 0.018 0.016 0.022
L2 distance of corr.  ($\downarrow$) 0.866 1.321 2.187 2.745 3.786 1.200 2.025 0.782 0.756 0.990
Detection score ($\downarrow$) 0.661 0.934 0.986 0.976 0.769 0.936 0.877 0.796 0.768 0.783
JSD ($\downarrow$) 0.055 0.011 0.114 0.152 0.051 0.038 0.048 0.015 0.016 0.018
WD ($\downarrow$) 0.004 0.011 0.023 0.025 0.061 0.022 0.016 0.010 0.007 0.009
DCR (abs. diff. to test; $\downarrow$) 1.278 1.588 3.336 1.621 0.568 1.000 2.593 0.796 0.806 0.758
@@
Table 12: L2 norm (incl. standard errors in subscripts) of the correlation matrix differences of real and synthetic train sets for seven benchmark models and for CDTD with three different noise schedules.
@@
SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
acsincome - $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.524_{\pm 0.010}$ $0.141_{\pm 0.003}$ $0.129_{\pm 0.003}$ $0.119_{\pm 0.002}$
adult $0.414_{\pm 0.016}$ $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.156_{\pm 0.006}$ $0.493_{\pm 0.009}$ $0.449_{\pm 0.011}$ $0.170_{\pm 0.007}$ $0.125_{\pm 0.009}$ $0.128_{\pm 0.010}$
bank $0.404_{\pm 0.015}$ $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.898_{\pm 0.025}$ $0.499_{\pm 0.021}$ $0.677_{\pm 0.015}$ $0.323_{\pm 0.008}$ $0.266_{\pm 0.011}$ $0.256_{\pm 0.015}$
beijing $0.081_{\pm 0.007}$ $0.133_{\pm 0.006}$ $1.445_{\pm 0.009}$ $1.642_{\pm 0.015}$ $1.133_{\pm 0.035}$ $0.363_{\pm 0.015}$ $0.096_{\pm 0.008}$ $0.075_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.071_{\pm 0.004}$
churn $0.264_{\pm 0.036}$ $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.327_{\pm 0.044}$ $0.746_{\pm 0.062}$ $0.509_{\pm 0.053}$ $0.302_{\pm 0.041}$ $0.289_{\pm 0.043}$ $0.282_{\pm 0.044}$
covertype - $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.044_{\pm 0.001}$ $1.029_{\pm 0.032}$ $3.958_{\pm 0.243}$ $2.359_{\pm 0.011}$ $2.275_{\pm 0.009}$ $2.710_{\pm 0.009}$
default $0.709_{\pm 0.048}$ $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $3.408_{\pm 0.105}$ $1.672_{\pm 0.061}$ $1.121_{\pm 0.042}$ $0.627_{\pm 0.068}$ $0.652_{\pm 0.102}$ $0.737_{\pm 0.033}$
diabetes $2.355_{\pm 0.026}$ $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $2.381_{\pm 0.026}$ $1.201_{\pm 0.020}$ $0.803_{\pm 0.032}$ $1.345_{\pm 0.016}$
lending $1.321_{\pm 0.063}$ $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.675_{\pm 0.015}$ $6.701_{\pm 0.034}$ $1.042_{\pm 0.075}$ $1.189_{\pm 0.040}$ $1.363_{\pm 0.097}$
news $1.684_{\pm 1.466}$ $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $15.985_{\pm 0.081}$ $4.874_{\pm 0.148}$ $4.990_{\pm 0.024}$ $1.925_{\pm 0.527}$ $2.035_{\pm 0.475}$ $3.395_{\pm 0.950}$
nmes $0.565_{\pm 0.047}$ $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.447_{\pm 0.031}$ $0.609_{\pm 0.032}$ $0.867_{\pm 0.046}$ $0.433_{\pm 0.025}$ $0.478_{\pm 0.083}$ $0.481_{\pm 0.058}$
@@
Table 13: Jensen-Shannon divergence (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
@@
SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
acsincome - $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.000}$ $0.022_{\pm 0.001}$
adult $0.064_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.034_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.020_{\pm 0.001}$ $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.016_{\pm 0.000}$
bank $0.039_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.029_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.054_{\pm 0.001}$ $0.010_{\pm 0.000}$ $0.009_{\pm 0.001}$ $0.012_{\pm 0.001}$
beijing $0.006_{\pm 0.002}$ $0.005_{\pm 0.002}$ $0.147_{\pm 0.003}$ $0.257_{\pm 0.001}$ $0.035_{\pm 0.003}$ $0.018_{\pm 0.004}$ $0.007_{\pm 0.002}$ $0.003_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.005_{\pm 0.001}$
churn $0.012_{\pm 0.004}$ $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.014_{\pm 0.004}$ $0.043_{\pm 0.001}$ $0.017_{\pm 0.002}$ $0.012_{\pm 0.003}$ $0.012_{\pm 0.002}$ $0.011_{\pm 0.002}$
covertype - $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.049_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.011_{\pm 0.000}$
default $0.042_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.027_{\pm 0.002}$ $0.073_{\pm 0.002}$ $0.082_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
diabetes $0.067_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.095_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.023_{\pm 0.000}$ $0.026_{\pm 0.000}$
lending $0.143_{\pm 0.001}$ $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.243_{\pm 0.002}$ $0.114_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.064_{\pm 0.002}$
news $0.063_{\pm 0.001}$ $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.046_{\pm 0.000}$ $0.012_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
nmes $0.060_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.028_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.026_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.013_{\pm 0.003}$
@@
Table 14: Wasserstein distance (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
@@
SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
acsincome - $0.007_{\pm 0.000}$ $0.037_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
adult $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
bank $0.002_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$
beijing $0.002_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.058_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
churn $0.006_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.013_{\pm 0.002}$ $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.006_{\pm 0.001}$
covertype - $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.015_{\pm 0.000}$ $0.012_{\pm 0.000}$
default $0.002_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
diabetes $0.004_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.042_{\pm 0.000}$
lending $0.006_{\pm 0.000}$ $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.425_{\pm 0.001}$ $0.050_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.010_{\pm 0.000}$ $0.011_{\pm 0.000}$
news $0.007_{\pm 0.000}$ $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.078_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.025_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
nmes $0.005_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$
@@
Table 15: Detection score (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.